OpenStudy (anonymous):

Which equation has the shape of the function f(x) = x^2 but moved to the left 2 units.

OpenStudy (anonymous):

Didn't read that correctly!

OpenStudy (anonymous):

f(X) = (x + 2)2 f(x) = (x - 2)2 f(x) = x2 + 2 f(x) = x2 - 2 which one of these would be the answer??